Could you please help me solve this equation? I keep coming up with answers other than the one that the book shows is right! Could you please show the steps you took to arrive at the answer? Thanks!!! (1/2 is one half - couldn't get it to show another way)
1/2(2n-5)=4n-1

Answer by nerdybill(7384)

1/2(2n-5)=4n-1
multiplying both sides by 2:
(2n-5)=8n-2
2n-5=8n-2
subtracting 2n from both sides:
-5=6n-2
adding two to both sides:
-3=6n
dividing both sides by 6:
-3/6 = n
-1/2 = n